A box contains 15 transistors, 3 of which are defective. If 3 are selected at random. the probability that.
a. all are defective
b.none are defective

Answer by Alan3354(69443)

A box contains 15 transistors, 3 of which are defective. If 3 are selected at random. the probability that.
----
Assuming you don't put them back (w/o replacement):
a. all are defective
= (3/15)*(2/14)*(1/13)
= 1/455
=============
b. none are defective
= (12/15)*(11/14)*(10/13)
= 220/455
= 44/91